OpenStudy (anonymous):

Write an equation or an inequality to model the real-life situation. Can you explain how you would write it? Pi is the quotient of the circumference "C" and the diameter "d" of a circle.

OpenStudy (anonymous):

\[\pi \sqrt{(C)(d)}\]

OpenStudy (anonymous):

So quotient is the same as a radical?

OpenStudy (anonymous):

it should be \[\pi=\sqrt{(C)(d)}\]

OpenStudy (anonymous):

sorry about that, I messed it up a tad bit. I'm too tired to think right now

OpenStudy (anonymous):

i'm pretty sure that's it though, 95.76% sure

OpenStudy (anonymous):

No, if the question is asking you to right "pi is the quotient of C and D" and not asking for your answer to be consistent with the relationships between pi, C and D, then the answer is: \[\pi=c/d\]

OpenStudy (anonymous):

*write, not right

OpenStudy (anonymous):

that sounds better than what i said, i didn't quite understand the question i guess

OpenStudy (anonymous):

And indeed, that's how you calculate pi: by dividing the circumference of a circle by its diameter

OpenStudy (anonymous):

Ok, thank you for explaining it to me. :)